OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 A. 2 B. 3 C. All real numbers D. No solution

OpenStudy (anonymous):

-2x-6=-2x-2-4 -2x-6=-2x-6 -2x+2x=-6x+6x 0=0 so D no solution

OpenStudy (anonymous):

oops -6+6 i meant